OpenStudy (anonymous):

what is the integral from ln6 to ln8 of coth

OpenStudy (anonymous):

sorry i typed it wrong

OpenStudy (anonymous):

\[\int\limits_{\ln6}^{\ln8}\coth4x\]

OpenStudy (anonymous):

i need it in ln not log

OpenStudy (zarkon):

it is in terms of \[\ln(\cdot)\] look that the small print on the link i provided "Log(x) is the natural logarithm"